40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11894     Entry time: Mon Dec 21 02:29:49 2015     In reply to: 11892     Reply to this: 11908
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX X RIN measurements 

I'll finish up the beat / frequency noise parts of the diagnosis tomorrow later, but I've done some investigation of the AUX X laser RIN. 

I placed a PDA255 at one of the rejected beams from the PBS on the downstream side of the IR faraday, making sure the power didn't saturate the PD. I measured the RIN on a SR785, and simultaneously looked at the signal on a 100MHz scope. 

The RIN has a very strong dependence on the laser diode current, and no noticable dependence on the crystal temperature or the presence of the PDH modulation / temperature control cables. Here are some traces, note that "nominal" current up until recently was 2.0A. 

When adjusting the diode current, a peak beings to appear in the tens of kHz, eventually noticible in the DC power trace on the scope. The point at which this occurs is not fixed.

At all times, I saw a strong intensity fluctuation at around 380-400kHz on the scope whose amplitude fluctuated a fair amount (at least 75mVrms over Vdc=6.5V, but would often be 2 or 3 times that).

I didn't look at the frequency noise while doing this, because the WiFi at the X end was too slow, I'll do more tomorrow in the daytime. 

Attachment 1: auxXRIN.pdf  72 kB  | Show | Show all
ELOG V3.1.3-